L Q AP MP

0 0 - -

1 8 8 8

2 22 11 14

3 60 20 38

4 96 ? 36

5 ? ? ?

6 132 22 ?



B1. If in the 5th worker the Average Product (AP) gets its maximum value, fill in the blanks of the table, where there is a question mark, presenting the appropriate calculations.


B2. Formulate the law of decreasing or disproportionate returns. In what quantity of the variable coefficient "labor" does the law of decreasing or disproportionate returns appear? Justify your answer.


B3. The variable cost (VC) of the business consists of labor costs and raw material costs. The wage of labor is w = 3.000 monetary units and is fixed for each worker.

The cost of raw materials (c) is fixed and equal for each unit of product produced.

Find the fixed cost (FC) of the business if at level of production of the 5th worker the marginal cost (MC) is 525 monetary units and the average total cost (ATC) is 700 monetary units.



1
Expert's answer
2021-11-25T12:44:46-0500

B1.



At the 4th worker, the average product would be 24, since:

Average product"=\frac{TP}{L}=\frac{96}{4}=24"

At the 6th worker, the marginal product would be 0, since, marginal product is zero at maximum value of total product.

Since marginal product at the 6th worker is zero, this implies that the total product at the 5th worker also 132.

The average product at the 5th worker is given as:

"\frac{132}{5}=26.4"

The value of marginal product in this case is also 26.4, because at maximum value of average product, marginal product is equal to marginal product.


B2.

The law of decreasing returns states that if one input in the production of a commodity is increased while all other inputs are held fixed, a point will eventually be reached at which additions of the input yield progressively smaller increases in output.

This law appears on the table when labor is 6.